# The Timescales of Embedded Star Formation as Observed in STARFORGE

### v1

@@ Abstract
Star formation occurs within dusty molecular clouds that are then disrupted by stellar feedback. However, the timing and physical mechanisms that govern the transition from deeply embedded to exposed stars remain uncertain. Using the STARFORGE simulations, we analyze the evolution of “embeddedness”, identifying what drives emergence. We find the transition from embedded to exposed is fast for individual stars, within 1.3 Myr after the star reaches its maximum mass. This rapid transition is dominated by massive stars, which accrete while remaining highly obscured until their feedback eventually balances, then overcomes, the local accretion. For these massive stars, their maximum mass is reached simultaneously with their emergence. Once these stars are revealed, their localized, pre-supernova feedback then impacts the cloud, driving gas clearance. Because massive stars dominate the luminosity, their fast, local evolution dominates the light emergence from the dust. We calculate the dependence of these processes on the mass of the cloud and find that emergence always depends on when massive stars form, which scales with the cloud’s free-fall time. We also examine the evolution of dust emission and H luminosity. We find that dust dominates the luminosity for roughly 2 Myrs before stellar luminosity becomes more luminous. These results suggest that deeply embedded star-forming clusters tend to be rare compared to those partially exposed. Thus, because the initial embedding of the most luminous stars is highly local, the emergence of stars is a faster, earlier, more local event than the overall disruption of the cloud by gas expulsion.
1 Introduction
@@
Efforts to constrain embedded cluster lifetimes have historically relied on associating young stellar clusters with their parent molecular clouds. Early studies found that clusters younger than Myr were often still embedded, while older ones were not, suggesting gas is typically cleared on Myr timescales (e.g., Leisawitz et al., 1989; Lada & Lada, 2003). More recent work extended this method to external galaxies using HST to detect young clusters and using ALMA observations to trace molecular gas or H to trace ionized gas (e.g., Whitmore et al., 2011; Grasha et al., 2018; Hannon et al., 2022). These larger samples often yield shorter inferred timescales of 1–2 Myr across diverse galactic environments.
@@
While there is general consensus that the de-embedding process takes less than 5 Myr (Portegies Zwart et al., 2010; Schinnerer & Leroy, 2024), the process is complex. Importantly, the interpretation of this timescale varies depending on the observable—whether it is CO, dust, PAH, or H emission. Different tracers probe different phases of feedback and dispersal (Kruijssen et al., 2018), and different techniques emphasize different parts of the sequence. Moreover, mounting evidence suggests that the embedded lifetime may not be fixed, but instead varies with the mass and environment of the parent cloud (McQuaid et al., 2024; Leroy et al., 2025; Knutas et al., 2025).
@@
This paper is structured as follows. We first introduce the STARFORGE simulations in Section 2. We then describe our methodology for measuring stellar properties (Section 3.1), and extinction (Section 3.2), and present the evolution of embeddedness in Sections 4.1, 4.3. We then present different wavelength observations of the simulation in Section 5, highlighting timescales in Section 5.4. We then discuss the implications of these measurements in Section 6.1.
2 STARFORGE Simulations
@@
In STARFORGE, a gas cell forms a sink particle once it satisfies a list of criteria designed to identify the centers of runaway collapse (see Grudić et al. 2021). The newly formed sink particle can then accrete gas based on the criteria of Bate et al. (1995), which include the gas gravitational boundedness, its proximity to the sink particle, angular momentum, and density.
Each sink particle contains a model star that accretes continuously from the sink’s internal mass reservoir, evolving properties such as the radius according to a semi-analytic protostellar evolution model Offner et al. (2009). Each stars’ winds are determined by the luminosity-dependent fitting formula of Grudić et al. (2021, 2022), where the emergent spectra are approximated as a blackbody with Teff = 5780K(L∗/L⊙)1/4(R∗/R⊙)-1/2. Stars may continue to accrete after reaching the main sequence, which we discuss further in Section 3.1.
In this work, we primarily focus on a fiducial simulation111named: M2e4_R10, which begins with a molecular cloud of uniform density. The cloud has a mass of 2000 M⊙, a radius of 10pc, in a 100pc periodic box filled with an ambient medium 1000 less dense than the original cloud. The flow is initialized as a Gaussian random velocity field with a power spectrum and scaled to obtain an initial virial parameter of 2, such that the cloud is marginally bound. The initial magnetic field is uniform and scaled so that the magnetic pressure is 10% of the turbulent ram pressure (Miville-Deschênes et al., 2017). The mass resolution is 10-3 M⊙ for gas cells, and 10-4 M⊙ for protostellar jets and stellar wind gas cells. In section 4.4, we also analyze a second simulation222named: M2e5_R30, which has a mass of 20000 M⊙, and a radius of 30pc but the same governing physics as the fiducial model.
3 Quantifying the Properties of Embedded Stars in STARFORGE
@@
3.1 Quantifying Stellar Ages
@@
We demonstrate these effects and show the resulting age spread in Figure 1. In the upper panel, we compare the simulation time when each protostar begins it’s protostellar phase (on the X-axis) to the time when the star reaches its final mass (on the Y-axis), color-coded by the final mass. The histogram in the upper panel shows the age distribution for the start of the protostellar phase, for all stars (red) and for the subset of more massive stars (; blue) 333We explored further binning the stars, but found that all bins above 1 M⊙ behaved the same.. The bottom panel shows the total duration of the protostars’ mass-accretion stage for the same mass bins.
Figure 1 clearly demonstrates the extended nature of star formation in the cloud. Protostars begin to form within 2 Myrs of the start of the simulation, and while the formation peaks at Myr, new protostars continue to form until well after 10 Myr, such that over 40% of stars form after 6.5 Myr, and the rms in the stellar formation times is 6.7 Myr.
In addition to the scatter in proto-star formation times, the accretion time to reach the final mass can be significant, particularly for the stars more massive than 1. These stars may take upwards of 3 Myr to reach their final masses (Grudić et al., 2022), depending on the availability of the gas supply. At early times, gas is ample, leading to longer accretion times for early-forming stars, on average.
We note, however, the accretion time does not decrease monotonically with time. This variation is driven by substructure in the cloud, such that the gas supply varies as a function of position. Some parts of the cloud may collapse at later times, in regions with less local feedback and thus larger local gas reservoirs. These local gas pockets can fuel on-going accretion, explaining the second period of high accretion between 7-10Myr.
3.1.1 Stellar Age Definition
@@
3.1.2 Exploring Stellar Age and Location
As the system dynamically evolves throughout the simulation, the location of both the stars, and the gas are constantly changing. As discussed in Section 3.1, we see a large age spread as different sub-clumps within the cloud collapse and form stars. In this section, we quantify the location of the stars with respect to their age.
The clumpy nature of the stars and gas is demonstrated in the interactive Figure 2. Shown in red are the 3 dimensional locations of stars, colored by their age (as defined in Section 3.1), where older stars are whiter, and the youngest stars still undergoing accretion are darker. We can clearly see that the youngest stars are tightly clustered in the bottom right corner of the still image. Readers are encouraged to rotate around simulation space and view the location of stars from several angles.
In addition to the stellar locations, the gray contours show gas density. At 8.8 Myr, we see small pockets of gas on the outskirts of the stellar association. Importantly, the location of the youngest stars directly corresponds to where the gas is the densest. This correlation between pockets of gas and newly formed stars demonstrates the sub-clump nature of the star formation in the model, and explains the large spread in stellar ages discussed in Section 3.1, where parts of the cloud collapse at different times.
One way to describe this correlation quantitatively is to measure the median distance to all other stars. This stellar density also can be compared between frames of the interactive Figure 2. Stars that form before 6.5 Myr are tightly clustered in physical space, with a median distance to other stars of 3.5pc . However, when we consider stars that form between 6.5 Myr, and 10 Myr, the median distance to other stars almost doubles to 6.8pc . This latter increased distance to other stars is consistent with a more-distributed mode of star formation that ensues when the already expanding cloud also begins disruption.
As the system evolves, the stars form in meaningfully different locations within the simulation box. Initially, the cloud just covers the 10pc and then collapses, which leads to stars forming centrally. However at later times, the cloud is expanding, and formation is occurring on the outskirts. Therefore the median formation location for stars that form before 6.5 Myr is 4.1 pc away from the median location of formation for stars after 6.5 Myr. This distance is also demonstrated in Figure 2 where the youngest stars are noticeably offset from the older stars.
3.2 Quantifying Extinction
@@
In Figure 3, the interquartile range is relatively constant in log space, which is consistent with the log-normal nature of most ISM density fields. We find that it typically spans around 00.5 in the log, independent of the median extinction, or the timestep.
@@
4 The Evolution of Embeddedness During Formation
@@
4.1 The Overall Evolution of Embeddedness
@@
We quantify this evolution in the bottom panel of Figure 4, where we show the fraction of stars that are deeply embedded ( mag; brown) or weakly embedded ( mag; red), as a function of time. These limits were chosen because in solar-metallicity, photo-dissociation models, mag is the approximate threshold for forming molecular hydrogen (Krumholz et al., 2009; Sternberg et al., 2014), where mag is a threshold within which carbon recombination takes place, and corresponds to HCN line emission (e.g., Hollenbach & Tielens, 1997; Kauffmann et al., 2017). The small vertical arrows connecting the top panels to the bottom plot indicate the times corresponding to the snapshot images. The times are given both in terms of the time since the start of the simulation (; bottom axis), and scaled by the free-fall time of the initial gas cloud (; top axis, calculated as Myr). Every point on the bottom plot includes all stars formed at or before , not just the stars forming at that particular timestep.
@@
4.2 Drivers of Embeddedness Evolution
@@
4.2.1 The Accounting of High-Mass Stars
@@
4.2.2 The Time Evolution of Drivers
@@
With the definition of “high mass” now settled, in Figure 6 we compare the time evolution of the embeddedness to the time evolution of a variety of possible drivers of stellar feedback to look for possible correlations. In both panels, we plot the time evolution of the embeddedness, tracked using the fraction of stars that are heavily embedded (; brown) or lightly embedded (;red), reproduced from Figure 4 and discussed in Section 4.1.
@@
In the left panel, we compare the change in the fraction of embedded stars to the fraction of stars which have begun their protostellar phase, for both stars with high final masses (M; dark blue), and all stars regardless of their final mass (black line). We count massive stars from the time accretion begins until the time they explode and normalize the curves by the maximum number of stars in each category at any point during the simulation. Note that this normalization leads to a downturn in the fraction of massive stars at 11Myr in Figure 6, when SNe begin.
@@
Comparing the number of stars to the embeddedness shows the expected impact of stellar feedback, namely that the stellar extinction (red and brown) drops systematically as more stars form. However, there is a surprisingly large amount of star formation before the embeddedness begins to fall sharply (4.7Myr, based on tracking , 6.0Myr, for ), with no previous clear impact on extinction. Quantitatively, more than 25% of all stars have already begun forming by this point. Perhaps more strikingly, close to 50% of eventually massive stars have begun forming before there is any significant drop in the typical embeddedness.
4.2.3 Only Considering Massive Stars Once they are Massive
@@
This unexpected result can be explained by considering how the stars are being counted. The left panel counts stars from the time they begin accretion, but it is clear from Figure 5 that even “eventually massive” stars can exist upwards of 2 Myr before they are meaningfully more massive than the other stars. As such, there is no appreciable difference between the blue and black curves, and no guarantee that any of the stars contributing to the curves currently have high mass, early in the simulation.
@@
In contrast, we see very different behavior in the right hand panel of Figure 6, where we choose to only consider massive stars once they surpass the 8 M⊙ threshold (i.e., the heavy lines in the left panel of Figure 5). With this choice, we are only counting massive stars once they are massive enough to have had either, significant mass accretion rates (i.e., significant protostellar feedback), or prolonged accretion events, such that they have large luminosities.
@@
With this different choice of how to define “massive”, we see a different correlation with the embeddedness that tracks dense gas. In contrast to the left panel, the decrease in extinction now appears to be more tightly related to the onset of massive stars (at 4-5Myr) rather than star formation in general (comparing the blue and grey curves). In other words, the drop in significant stellar embeddeding is driven by effects produced by stars as they reach a significant mass.
@@
While the observed correlation strongly suggests that massive stars are driving the drop in extinction, there are multiple possible mechanisms at work. We consider one of these mechanisms in the right panel of Figure 6, where we assess the role of radiation, which is dominated by the population of hot massive stars. We plot the time evolution of the bolometric stellar luminosity (green) and the number of ionizing photons (purple), each normalized by their maximum value throughout the simulation444For reference, the maximum number of ionizing photons is photons, and the maximum stellar luminosity is L⊙..
@@
As expected, both the luminosity and the number of ionizing photons increase along with the number of stars. Even more striking is the degree to which both quantities track the subset of massive ( M⊙) stars, rising in lockstep until 6Myr, which is roughly the time by which half of all stars (massive and otherwise) have formed, half the sight-lines are no longer highly embedded (brown curve), and the fraction of lightly embedded stars begins to decline (red curve). At later times, the total stellar luminosity and the number of ionizing photons grow even faster, presumably as a subset of stars reach masses higher than the 8 M⊙ threshold used as the mass limit for the blue curve.
@@
4.3 Tracking the Evolution of Formation Time and Embeddedness for Individual Stars

### v2

@@ Abstract
Star formation occurs within dusty molecular clouds that are then disrupted by stellar feedback. However, the timing and physical mechanisms that govern the transition from deeply embedded to exposed stars remain uncertain. Using the STARFORGE simulations, we analyze the evolution of “embeddedness”, identifying what drives emergence. We find the transition from embedded to exposed is fast for individual stars, within 1.3 Myr after the star reaches its maximum mass. This rapid transition is dominated by massive stars, which accrete while remaining highly obscured until their feedback eventually balances, then overcomes, the local accretion. For these massive stars, their maximum mass is reached simultaneously with their emergence. Once these stars are revealed, their localized, pre-supernova feedback then impacts the cloud, driving gas clearance. Because massive stars dominate the luminosity, their fast, local evolution dominates the light emergence from the dust. We calculate the dependence of these processes on the mass of the cloud and find that emergence always depends on when massive stars form, which scales with the cloud’s free-fall time. We also measure the evolution of dust and H luminosities, where for 2 Myr, these tracers outshine the emerging stellar continuum, reaching their peak when gas and dust remain tightly coupled to the massive stars. These results closely resemble observationally observed lifetimes, tying the observable dust and line emission directly to the same localized processes that drive stellar emergence, evidence that our simulated de-embedding physics is representative of real star-forming regions. Thus, because the initial embedding of the most luminous stars is highly local, the emergence of stars is a faster, earlier, more local event than the overall disruption of the cloud by gas expulsion.
I Introduction
@@
Efforts to constrain embedded cluster lifetimes have historically relied on associating young stellar clusters with their parent molecular clouds. Early studies found that clusters younger than Myr were often still embedded, while older ones were not, suggesting gas is typically cleared on Myr timescales (e.g., Leisawitz et al., 1989; Whitmore & Zhang, 2002; Lada & Lada, 2003). More recent work extended this method to external galaxies using HST to detect young clusters and using ALMA observations to trace molecular gas or H to trace ionized gas (e.g., Whitmore et al., 2011; Grasha et al., 2018; Hannon et al., 2022; Whitmore et al., 2025). These larger samples often yield shorter inferred timescales of 1–2 Myr across diverse galactic environments.
@@
While there is general consensus that the de-embedding process takes less than 5 Myr (Portegies Zwart et al., 2010; Schinnerer & Leroy, 2024), the process is complex. Importantly, the interpretation of this timescale varies depending on the observable—whether it is CO, dust, PAH, or H emission. Different tracers probe different phases of feedback and dispersal (Kruijssen et al., 2018), and different techniques emphasize different parts of the sequence (Leisawitz et al., 1989). Moreover, mounting evidence suggests that the embedded lifetime may not be fixed, but instead varies with the mass and environment of the parent cloud (McQuaid et al., 2024; Leroy et al., 2025; Knutas et al., 2025).
@@
This paper is structured as follows. We first introduce the STARFORGE simulations in Section II. We then describe our methodology for measuring stellar properties (Section III.1), and extinction (Section III.2), and present the evolution of embeddedness in Sections IV.1, IV.3. We then present different wavelength observations of the simulation in Section V, highlighting timescales in Section V.4. We then discuss the implications of these measurements in Section VI.1.
II STARFORGE Simulations
@@
In STARFORGE, a gas cell forms a sink particle once it satisfies a list of criteria designed to identify the centers of runaway collapse (see Grudić et al. 2021). The newly formed sink particle can then accrete gas based on the criteria of Bate et al. (1995), which include the gas gravitational boundedness, its proximity to the sink particle, angular momentum, and density. Once a particle has become a sink, its equation of motion includes only gravity and the back-reaction of local momentum conservation during accretion. Binary and higher-order multiple systems can form through the dynamics of distinct sinks (e.g., Guszejnov et al., 2023), but we do not resolve multiplicity for separations below au and each sink is treated as a single star for the purposes of this work.
Each sink hosts a single model protostar and an internal gas “reservoir”. Resolved accretion moves gas from the simulation domain into the sink by merging the discrete gas cells, and unresolved accretion then transfers mass from the reservoir onto the protostar, providing a continuous accretion rate in this regime (Grudić et al., 2021)111See Figure 4 in Grudić et al. (2021) for a complete picture of the sink, star, reservoir, and resulting feedback.. The protostar, and its properties such as luminosity and radius evolves through standard pre–collapse and protostellar phases to the main sequence (Offner et al., 2009). We approximate the emergent spectrum as a blackbody with . Protostars can continue to accrete after reaching the main sequence, which we discuss further in Section III.1.
Feedback is tied directly to this evolving protostar: protostellar outflows scale with the accretion rate, while radiative feedback and line–driven winds scale with the stellar mass and luminosity (Grudić et al., 2021, 2022). There is thus no single discrete “onset” of feedback; mechanical feedback is present from the moment a sink forms and radiative/wind feedback strengthen as massive stars grow. Self–gravity is solved on a finite, non–periodic domain whose size is 10 times the initial cloud radius, so the cloud evolves as an effectively isolated object rather than in a periodically replicated volume.
In this work, we primarily focus on a fiducial simulation222named: M2e4_R10, which begins with a molecular cloud of uniform density. The cloud has a mass of 2000 M⊙, a radius of 10pc, in a 100pc periodic box filled with an ambient medium 1000 less dense than the original cloud. The flow is initialized as a Gaussian random velocity field with a power spectrum and scaled to obtain an initial virial parameter of 2, such that the cloud is marginally bound. The initial magnetic field is uniform and scaled so that the magnetic pressure is 10% of the turbulent ram pressure (Miville-Deschênes et al., 2017). The mass resolution is 10-3 M⊙ for gas cells, and 10-4 M⊙ for protostellar jets and stellar wind gas cells. In section IV.4, we also analyze a second simulation333named: M2e5_R30, which has a mass of 20000 M⊙, and a radius of 30pc but the same governing physics as the fiducial model.
III Quantifying the Properties of Embedded Stars in STARFORGE
@@
III.1 Quantifying Stellar Ages
@@
We demonstrate these effects and show the resulting age spread in Figure 1. In the bottom panel, for each star we compare the simulation time when each protostar begins it’s protostellar phase (on the X-axis) to the time when the star reaches its final mass (on the Y-axis), color-coded by the final mass. The histogram in the upper panel shows the age distribution for the start of the protostellar phase, for all stars (red) and for the subset of more massive stars (; blue) 444We explored further binning the stars, but found that all bins above 1 M⊙ behaved the same.. The middle panel shows the total duration of the protostars’ mass-accretion stage for the same mass bins.
The density distribution histograms of stellar age, and range of points in the upper panel of Figure 1 clearly demonstrates the extended nature of star formation in the cloud. Protostars begin to form within 2 Myrs of the start of the simulation, and while the formation peaks at Myr, new protostars continue to form until well after 10 Myr, such that over 40% of stars form after 6.5 Myr, and the rms in the stellar formation times is 6.7 Myr.
In addition to the scatter in proto-star formation times, the accretion time to reach the final mass can be significant.In the bottom panel of Figure 1, for each point, this accretion time is the integrated distance between the star, and the black dashed line, while in the middle panel, this time is shown on the y-axis. Of particular interest is the accretion time for stars move massive than 1, which may take upwards of 3 Myr to reach their final masses (Grudić et al., 2022), depending on the availability of the gas supply. At early times, gas is ample, leading to longer accretion times for early-forming stars, on average.
We note, however, the accretion time does not decrease monotonically with time. This variation is driven by substructure in the cloud, such that the gas supply varies as a function of position. Some parts of the cloud may collapse at later times, in regions with less local feedback and thus larger local gas reservoirs. These local gas pockets can fuel on-going accretion, explaining the second period where accretion rates for stars more massive than 1 M⊙ rise above 1 Myr between 7-10Myr.
III.1.1 Stellar Age Definition
@@
III.1.2 Exploring Stellar Age and Location
As the system dynamically evolves throughout the simulation, the location of both the stars, and the gas are constantly changing. As discussed in Section III.1, we see a large age spread as different sub-clumps within the cloud collapse and form stars. In this section, we quantify the location of the stars with respect to their age.
Throughout the simulation, the stars form in meaningfully different locations. Initially, the cloud spans 20pc and then collapses, which leads to stars forming centrally. However at later times, the cloud expands and formation occurs on the outskirts. Therefore, the median formation location for stars that form before 6.5 Myr is 4.1 pc away from the median location of formation for stars after 6.5 Myr. This distance is also demonstrated in Figure 2 where the youngest stars are noticeably offset from the older stars.
The clumpy nature of the stars and gas is demonstrated in the interactive Figure 2. Shown in red are the 3 dimensional locations of stars, colored by their age (as defined in Section III.1), where older stars are whiter, and the youngest stars still undergoing accretion are darker. We can clearly see that the youngest stars are tightly clustered in the bottom right corner of the still image. Readers are encouraged to rotate around simulation space and view the location of stars from several angles.
In addition to the stellar locations, the gray contours show gas density. At 8.8 Myr, we see small pockets of gas on the outskirts of the stellar association. Importantly, the location of the youngest stars directly corresponds to where the gas is the densest. This correlation between pockets of gas and newly formed stars demonstrates the sub-clump nature of the star formation in the model, and explains the large spread in stellar ages discussed in Section III.1, where parts of the cloud collapse at different times.
One way to describe this correlation quantitatively is to measure the median distance to all other stars. This stellar density also can be compared between frames of the interactive Figure 2. Stars that form before 6.5 Myr are tightly clustered in physical space, with a median distance to other stars of 3.5pc . However, when we consider stars that form between 6.5 Myr, and 10 Myr, the median distance to other stars almost doubles to 6.8pc . This latter increased distance to other stars is consistent with a more-distributed mode of star formation that ensues when the already expanding cloud also begins disruption. This increased distance is also clearly seen in the interactive Figure 2 by adjusting the slider between times; the stars become less tightly clustered as time progresses.
III.2 Quantifying Extinction
@@
In the left most panels of Figure 3, the interquartile range is relatively constant in log space, which is consistent with the log-normal nature of most ISM density fields. We find that it typically spans around 00.5 in the log, independent of the median extinction, or the timestep.
@@
IV The Evolution of Embeddedness During Formation
@@
IV.1 The Overall Evolution of Embeddedness
@@
We quantify this evolution in the bottom panel of Figure 4, where we show the fraction of stars that are deeply embedded ( mag; brown) or weakly embedded ( mag; red), as a function of time. These limits were chosen because in solar-metallicity, photo-dissociation models, mag is the approximate threshold for forming molecular hydrogen (Franco & Cox, 1986; Krumholz et al., 2009; Sternberg et al., 2014), where mag is a threshold within which carbon recombination takes place, and corresponds to HCN line emission (e.g., Hollenbach & Tielens, 1997; Kauffmann et al., 2017). The small vertical arrows connecting the top panels to the bottom plot indicate the times corresponding to the snapshot images. The times are given both in terms of the time since the start of the simulation (; bottom axis), and scaled by the free-fall time of the initial gas cloud (; top axis, calculated as Myr). Every point on the bottom plot includes all stars formed at or before , not just the stars forming at that particular timestep.
@@
IV.2 Drivers of Embeddedness Evolution
@@
IV.2.1 The Accounting of High-Mass Stars
@@
IV.2.2 The Time Evolution of Drivers
@@
With the definition of “high mass” now settled, in Figure 6 we compare the time evolution of the embeddedness to the time evolution of a variety of possible drivers of stellar feedback to look for possible correlations. In both panels, we plot the time evolution of the embeddedness, tracked using the fraction of stars that are heavily embedded (; brown) or lightly embedded (;red), reproduced from Figure 4 and discussed in Section IV.1.
@@
In the top panel, we compare the change in the fraction of embedded stars to the fraction of stars which have begun their protostellar phase, for both stars with high final masses (M; dark blue), and all stars regardless of their final mass (black line). We count massive stars from the time accretion begins until the time they explode and normalize the curves by the maximum number of stars in each category at any point during the simulation. Note that this normalization leads to a downturn in the fraction of massive stars at 11Myr in Figure 6, when SNe begin.
@@
Comparing the number of stars to the embeddedness shows the expected impact of stellar feedback, namely that the stellar extinction (highly transparent red and brown lines) drops systematically as more stars form. However, a surprisingly large amount of star formation occurs before we see any clear decline in extinction. Specifically, the remains above 90% until 4.7Myr, and remains above 90% until 6.0Myr, when both begin to decline more rapidly. Quantitatively, at 4.7Myr, only 25% of all stars have already begun forming. Perhaps more strikingly, close to 50% of eventually massive stars have begun forming by 6.0Myr, when there is a significant drop in the typical embeddedness.
IV.2.3 Only Considering Massive Stars Once they are Massive
@@
This unexpected result can be explained by considering how the stars are being counted. The top panel counts stars from the time they begin accretion, but it is clear from Figure 5 that even “eventually massive” stars can exist upwards of 2 Myr before they are meaningfully more massive than the other stars. As such, there is no appreciable difference between the blue and black curves in the top panel of Figure 6, and no guarantee that any of the stars contributing to the curves currently have high mass, early in the simulation.
@@
In contrast, we see very different behavior in the bottom panel of Figure 6, where we choose to only consider massive stars once they surpass the 8 M⊙ threshold (i.e., above the dashed line in the left panel of Figure 5). With this choice, we are only counting massive stars once they are massive enough to have had either, significant mass accretion rates (i.e., significant protostellar feedback), or prolonged accretion events, such that they have large luminosities.
@@
With this different choice of how to define “massive”, we see a different correlation with the embeddedness that tracks dense gas. In contrast to the top panel, the decrease in extinction now appears to be more tightly related to the onset of massive stars (at 4-5Myr) rather than star formation in general (comparing the blue and grey curves). In other words, the drop in significant stellar embeddeding is driven by effects produced by stars as they reach a significant mass.
@@
While the observed correlation strongly suggests that massive stars are driving the drop in extinction, there are multiple possible mechanisms at work. We consider one of these mechanisms in the bottom panel of Figure 6, where we assess the role of radiation, which is dominated by the population of hot massive stars. We plot the time evolution of the bolometric stellar luminosity (green) and the number of ionizing photons (purple), each normalized by their maximum value throughout the simulation555For reference, the maximum number of ionizing photons is photons, and the maximum stellar luminosity is L⊙..
@@
As expected, both the luminosity and the number of ionizing photons increase along with the number of stars. Even more striking is the degree to which both quantities track the subset of massive ( M⊙) stars, rising in lockstep until 6Myr, which is roughly the time by which half of all stars (massive and otherwise) have formed, half the sight-lines are no longer highly embedded (transparent brown curve), and the fraction of lightly embedded stars begins to decline (transparent red curve). At later times, the total stellar luminosity and the number of ionizing photons grow even faster, presumably as a subset of stars reach masses higher than the 8 M⊙ threshold used as the mass limit for the blue curve.
@@
IV.3 Tracking the Evolution of Formation Time and Embeddedness for Individual Stars